# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract.
Secure collaborative analytics (SCA) enable the processing of analytical SQL queries across multiple owners’ data, even when direct data sharing is not feasible. Although essential for strong privacy, the large overhead from data-oblivious primitives in traditional SCA has hindered its practical adoption. Recent SCA variants that permit controlled leakages under differential privacy (DP) show a better balance between privacy and efficiency. However, they still face significant challenges, such as potentially unbounded privacy loss, suboptimal query planning, and lossy processing.
To address these challenges, we introduce SPECIAL, the first SCA system that simultaneously ensures bounded privacy loss, advanced query planning, and lossless processing. SPECIAL employs a novel synopsis-assisted secure processing model, where a one-time privacy cost is spent to acquire private synopses (table statistics) from owner data. These synopses then allow SPECIAL to estimate (compaction) sizes for secure operations (e.g., filter, join) and index encrypted data without extra privacy loss. Crucially, these estimates and indexes can be prepared before runtime, thereby facilitating efficient query planning and accurate cost estimations. Moreover, by using one-sided noise mechanisms and private upper bound techniques, SPECIAL ensures strict lossless processing for complex queries (e.g., multi-join). Through a comprehensive benchmark, we show that SPECIAL significantly outperforms cutting-edge SCAs, with up to 80×80$\times$80 × faster query times and over 900×900$\times$900 × smaller memory for complex queries. Moreover, it also achieves up to an 89×89$\times$89 × reduction in privacy loss under continual processing.
@@ 1. Introduction
While MPC can effectively conceal data values (Yao, 1986), its security guarantees do not immediately extend to the protection of execution transcripts. Consequently, data-dependent processing patterns such as memory traces and read/write volumes can still reveal critical information, risking privacy breaches (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016) even when the core data remains encrypted. To ensure strong privacy, modern SCA systems also utilize data-oblivious primitives that exhaustively pad query processing complexities to a worst-case, data-independent upper bound (Bater et al., 2017; Liagouris et al., 2023; Poddar et al., 2021). However, such stringent protections can largely reduce system efficiency and hinder the generalization of conventional optimization techniques to SCA, which are typically data-dependent (Liagouris et al., 2023). To address this, recent efforts (Qin et al., 2022; Wang et al., 2021, 2022; Bater et al., 2018) have introduced differentially private SCA (DPSCA). This approach allows controlled information leakage under differential privacy (Dwork et al., 2014) to mitigate constant worst-case overhead. For instance, systems under this model can dynamically compact intermediate query sizes to a noisy estimate close to actual sizes, avoiding exhaustive padding. As such, queries under DPSCA experience largely boosted efficiencies (e.g., up to $10^{5}\times$ faster (Wang et al., 2022)) compared to their “no leakage” counterparts. Despite these substantial performance gains, existing DPSCAs still face critical limitations that impede their practical uses, as elaborated below:
@@
$\bullet$ L-1. Unbounded privacy loss. Most DPSCA systems utilize a per-operator privacy expenditure model (Bater et al., 2018; Wang et al., 2021; Qin et al., 2022; Chu et al., 2021; Chan et al., 2022; Wang et al., 2022), meaning each query operator (e.g., join, filter) independently consumes a portion of the privacy budget. This approach can lead to either unbounded privacy loss or the forced cessation of query responses upon budget exhaustion. To mitigate this, some studies (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]) propose private, locality-sensitive grouping, incurring a one-time privacy cost to pre-group data based on specific attributes. Subsequent queries on those attributes can be directly applied to a smaller subset and need no additional privacy budget. However, this method only supports simple queries (e.g., point and range); complex queries like joins still suffer from unbounded privacy loss.
@@
$\bullet$ L-2. Suboptimal execution plan. Conventional query planners can pre-estimate sizes for equivalent plans of a given query and select the most efficient plan with minimized intermediate sizes before execution (Silberschatz et al., 2011; Blasgen et al., 1981). In contrast, SCA systems lack this capability, and even DPSCA designs (Qin et al., 2022; Wang et al., 2022; Zhang et al., 2023; Wagh et al., 2021; Bater et al., 2018, 2020) can only reactively determine plan sizes during runtime. This inherent limitation often forces existing systems to settle for less efficient query plans, such as suboptimal join orders, which lead to significantly inflated intermediate sizes (§ 7.2) and substantially hinder performance.
@@
$\bullet$ L-3. Lossy processing. Noise from randomized mechanisms in DPSCA also introduces a unique accuracy issue (e.g., losing tuples after compaction), and unfortunately, no existing design can provide deterministic accuracy guarantees for complex queries (Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Zhang et al., 2023; Groce et al., 2019). Furthermore, stronger privacy settings can further increase noise variance, which amplifies errors, significantly impacting the utility of SCA systems.
1.1. Overview of SPECIAL.
@@
In this work, we introduce SPECIAL, an innovative SCA system designed to address all these challenges simultaneously through a new paradigm called synopsis-assisted secure processing. At its core, SPECIAL incurs a one-time privacy cost to gather DP synopses (statistics of base tables) from owners’ data. These synopses are then used to compact query intermediate results, index encrypted data, and enhance SCA query planning. Notably, SPECIAL is the first system to provide all of the following benefits: (1) Bounded privacy—it manages complex queries, such as multi joins, within strict privacy limits; (2) Advanced query planning—it builds an advanced SCA planner that can exploit plan sizes before runtime; and (3) Lossless processing—it ensures exact results with no data omissions. An overview of SPECIAL’s architecture is shown in Figure 1.
@@
SPECIAL operates under a standard server-aided MPC model (Kamara et al., 2011) with three key participants: data owners, at least two SPECIAL servers, and a vetted analyst. The process starts with data owners securely outsourcing their data (e.g. using secret sharing (Kamara et al., 2011; Wang et al., 2021; Zhang et al., 2023; Wang et al., 2022)), and privately release a set of DP synopses (table statistics) to the servers. Once the data and synopses are ready, analysts can then submit queries on the outsourced data. To process queries, SPECIAL uses synopses to privately estimate tight result compaction bounds or build private indexes for faster access, thus avoiding exhaustive padding. Crucially, this information is exploit before runtime, so that SPECIAL can pre-identify an optimal secure execution plan that minimizes intermediate sizes and overall costs. Finally, the optimized plan is executed via MPC across the secret-shared data, and the results are securely returned to the analyst.
In the SPECIAL prototype, we adopt the same privacy model used in existing DPSCA designs, assuming a computationally bounded, “honest-but-curious” adversary (Goldreich, 2004) who can compromise up to $n-1$ out of $n$ owners and $m-1$ out of $m$ servers. SPECIAL ensures that such an adversary can only learn a limited leakage profile $\mathsf{Lkg}$ about honest parties’ data, which strictly adheres to DP constraints. Note that, SPECIAL’s core components, including synopses management and query planning do not leverage raw data at all. Thus, any of the servers, alone or jointly, can manage these tasks. Also, once analysts are authorized by data owners, they are considered trusted, and additional randomization of query results is not required. It is worth mentioning that SPECIAL’s corruption model (e.g. all except one) is entirely determined by the underlying MPC, and SPECIAL’s design does not modify them at all. Hence, one may also opt to a weaker corruption model, such as a supermajority of owners and servers must remain uncorrupted (Liagouris et al., 2023; Tan et al., 2021), to enhance efficiency.
@@ 1.2. Unique challenges and key contributions
$\bullet$ C-1. How to select proper synopses? Even a single relation can have numerous options for generating synopses, such as using different attributes or their combinations. However, improper selection can lead to large errors (e.g. using too many synopses or leverage high-dimensional attributes (Zhang et al., 2021)), or reduced functionalities (e.g., using only simple attributes (Zhang et al., 2023)). Hence, a key challenge is selecting a limited set of DP synopses to optimize the privacy budget for complex query processing. Our approach is informed by two observations: (i) secure joins are resource-intensive and need prioritized acceleration, and (ii) synopses for common filtering predicates are vital as they allow pre-built indexes on base relations for fast access. Consequently, we propose a focused strategy (§ 4) that targets low-dimensional (1D and 2D) attributes frequently involved in joins and filters within a representative workload.
@@
$\bullet$ C-2. How to enforce lossless processing? Private synopses do not immediately implies lossless guarantees. Thus, a second challenge is designing practical approaches to achieve lossless results without violating privacy goals. To address this, we employ one-sided DP noise (either strictly positive or negative, § 4) for generating synopses and design novel primitives (§ 5) based on them to pessimistically estimate filter cardinalities and intervals of index structures. This enable us to efficiently discard unnecessary records during processing without breaking privacy and lossless guarantees. To ensure lossless processing of complex joins, we extend upon cutting-edge join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes using available synopses. To our knowledge, this is the first study to support private join upper bound estimation.
@@
$\bullet$ C-3. How DP synopses can empower efficient query processing? The use of DP synopses in SCA is largely underexplored, leaving a knowledge gap regarding their potential to enhance query efficiency. To navigate this potential, we explore various use of synopses in accelerating secure processing including private indexes SPEidx (§ 5.2), and compacted oblivious operations SPEop (§ 5.3). We then develop a Selinger-style query planner (§ 6) that supports both standard secure operators and SPEop. Employing a customized cost model and tailored heuristics, this planner efficiently generates optimal execution plans for queries using a mix of standard and SPEop operators. This in essence, also address the limitations of sub-optimal execution plans inherent in standard DPSCA.
@@
$\bullet$ C-4. How to systematically evaluate SPECIAL? A major challenge in SCA design evaluation is the absence of open-source benchmarks. We address this by initiating an open-source evaluation platform accessible to the public, utilizing public Financial data (Rel, [n.d.]) (1.1M rows and 59 columns), and designing eight test queries ranging from simple linear to complex 5-way joins. We evaluated our prototype SPECIAL alongside the state-of-the-art DPSCA system, Shrinkwrap (Bater et al., 2018), and conventional SCA, SMCQL (Bater et al., 2017). Results indicate that SPECIAL outperforms Shrinkwrap, reducing query latency by up to 80.3×80.3$\times$80.3 ×, and SMCQL, with at least a 114×114$\times$114 × reduction in query latency. Additionally, SPECIAL improves memory efficiency in complex join processing by more than 900×900$\times$900 × compared to both systems. Moreover, scaling experiments show that SPECIAL can effectively scale linear and binary joins up to 8×8$\times$8 × (8.8M rows) and 5-way joins up to 4×4$\times$4 × (4.4M rows) workloads. All benchmarks, including our prototype implementation, are open-sourced and available at (Lovingmage, 2024).
The remainder of the paper is structured as follows: We begin with essential background in § 2 and a formal definition of our privacy model in § 3 before delving into design details. § 4 discusses the SPECIAL catalog design, including synopses selection and generation. § 5 details novel synopses-assisted secure database operators, and § 6 describes our planner design. Benchmark evaluations are presented in § 7, followed related works and conclusions.
@@ 2. Background
General notations. We consider the logical database $\mathcal{D}$ to contain multiple private relations $\{D_{1},D_{2},...\}$, where each relation $D_{i}$ is belongs owned by a specific party $P_{i}$, where $D\in\mathcal{D}$ is owned by a specific owner and has a set of attributes as $attr(D)$. The domain for attribute $A\in attr(D)$ is given by $dom(A)$, and the combined domain for a collection of attributes $\mathbf{A}=\{A_{1},A_{2},...\}\subseteq attr(D)$ is denoted as $dom(\mathbf{A})=\prod_{A\in\mathbf{A}}dom(A)$. For a tuple $t\in D$, and $\mathbf{A}\subseteq attr(D)$, we denote $t.\mathbf{A}$ as the attribute value of $\mathbf{A}$ in $t$
@@
Frequency (count). Given $D$, and $\mathbf{A}\subseteq attr(D)$, and $\mathbf{v}\in dom(\mathbf{A})$, the frequency (count) of $\mathbf{v}$ in $D$ is the total number of tuples $t\in D$ with $t.\mathbf{A}=\mathbf{v}$. In addition, the max frequency moments (MF) of $\mathbf{A}$ is defined as ${\mathtt{mf}}(\mathbf{A},D)=\max_{\mathbf{v}\in dom(\mathbf{A})}|\{t\in D\mid t.\mathbf{A}=\mathbf{v}\}|$.
@@
Histograms. Given $D$, and $\mathbf{A}\subseteq attr(D)$, the (equal-width) histogram $\mathbf{h}(\mathbf{A},D)=(c_{1},c_{2},...,c_{m})$ is a list of counts for the attribute values in $\mathbf{A}$. Specifically, $\mathbf{h}$ partitions [^1] We use a condensed notation that omits arguments when $\mathbf{A}$ and $D$ are obvious. $dom(\mathbf{A})$ into $m$ “equal-sized” domain intervals $(B_{1},...,B_{m})$, and each count $c_{i}\in\mathbf{h}$ corresponds to the total number of $t\in D$ such that $t.\mathbf{A}$ fall within $B_{i}$.
Multi-party secure computation (MPC). MPC (Yao, 1986; Goldreich, 2009; Micali et al., 1987; Ben-Or et al., 2019) is a cryptographic techniques to allow a group of participants $P_{1},P_{2},...$ to jointly compute a function $f(x_{1},x_{2},...)$ over their respective private input $x_{i}$. MPC ensures that no unauthorized information is revealed to any involved party, excepted the desired output of function $f$, emulating a computation as if performed by a trusted third party. Earlier MPC typically require all parties to participate in intensive computations. Recent schemes (Kamara et al., 2011; Scholl et al., 2017; Mohassel and Zhang, 2017) allow these computations to be offloaded to multiple powerful servers without sacrificing original security guarantees, known as the server-aided MPC. In this approach, each party, say $P_{i}$, secretly shares their inputs, $x_{i}$ to servers. These servers then jointly evaluate an MPC that functionally replicates the logic of first reconstructing the secrets $x_{1},x_{2},\ldots$ and then evaluating $f(x_{1},x_{2},\ldots)$. Depending on designs, the result may be directly returned to the parties or re-shared among the servers for later retrieval. By default, SPECIAL considers server-aided MPC with two non-colluding servers.
Secret sharing and secure array. SPECIAL uses the $2$-out-of-$2$ boolean secret share (Araki et al., 2016) over ring $\mathbb{Z}_{2^{32}}$ for securely outsourcing owners’ data and storing query execution results. Specifically, each data, $x$, is divided into two shares: $x1,x2$ that are uniformly distributed over the ring $\mathbb{Z}_{2^{32}}$ such that $x=x1\oplus x2$. Each server $S_{i}$ receives one secret shares, $s_{i}$, where $i\in\{0,1\}$. By retrieving shares from any two servers, an authorized party can successfully reconstruct the value of $x$. However, a single server alone learns nothing about $x$. For clarity and to abstract out the lower-level details, we leverage a logically unified data structure, namely the secure array (Wang et al., 2022; Bater et al., 2018), denoted as $\langle\mathbf{x}\rangle=\left(\langle x_{1}\rangle,\langle x_{2}\rangle,...\right)$, which is a collection of secret-shared relational tuples.
Oblivious (relational) operators. Oblivious operators are specific MPC protocols, applied over secure arrays, that implement database operator functionalities such as filter, join, and aggregation. Specifically, an oblivious operator $\langle\mathbf{o}\rangle\leftarrow\mathsf{\textbf{op}}(\langle\mathbf{a}\rangle,\langle\mathbf{b}\rangle,...)$ accepts one or more secure arrays as inputs, performs database operations over the inputs, and then writes the result to an output secure array. Moreover, the oblivious property stipulates that absolutely no memory access pattern information is disclosed (Chang et al., 2022; Ramachandran and Shi, 2021; Liagouris et al., 2023) (or in other words, the memory access is fully independent from input data) during the protocol’s execution. Since revealing output size can also enable numerous reconstruction attacks, recent designs of oblivious operators also consider hiding output volumes, i.e., by padding the outputs to their maximum sizes. Thus, unless otherwise stated, we assume that all the subsequent oblivious operators mentioned have considered volume-hiding.
@@
Differential privacy (Dwork et al., 2014). DP ensures that modifying a single input data tuple to a mechanism produces only a negligible change in its output. To elaborate, consider $D$ and $D^{\prime}$ as two databases differing by just one tuple—termed as neighboring databases. Then:
Definition 2.1 ($(\epsilon,\delta)$-differential privacy).

### v2

@@ Abstract.
Secure collaborative analytics (SCA) enables the processing of analytical SQL queries across data from multiple owners, even when direct data sharing is not possible. While traditional SCA provides strong privacy through data-oblivious methods, the significant overhead has limited its practical use. Recent SCA variants that allow controlled leakages under differential privacy (DP) strike balance between privacy and efficiency but still face challenges like unbounded privacy loss, costly execution plan, and lossy processing.
To address these challenges, we introduce SPECIAL, the first SCA system that simultaneously ensures bounded privacy loss, advanced query planning, and lossless processing. SPECIAL employs a novel synopsis-assisted secure processing model, where a one-time privacy cost is used to generate private synopses from owner data. These synopses enable SPECIAL to estimate compaction sizes for secure operations (e.g., filter, join) and index encrypted data without additional privacy loss. These estimates and indexes can be prepared before runtime, enabling efficient query planning and accurate cost estimations. By leveraging one-sided noise mechanisms and private upper bound techniques, SPECIAL guarantees lossless processing for complex queries (e.g., multi-join). Our comprehensive benchmarks demonstrate that SPECIAL outperforms state-of-the-art SCAs, with up to $80\times$ faster query times, $900\times$ smaller memory usage for complex queries, and up to $89\times$ reduced privacy loss in continual processing.
PVLDB Reference Format:
PVLDB, 18(4): 1035 - 1048, 2024.
doi:10.14778/3717755.3717764 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 4 ISSN 2150-8097.
doi:10.14778/3717755.3717764
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/lovingmage/caplan.
@@ 1. Introduction
While MPC can effectively conceal data values (Yao, 1986), its security guarantees do not immediately extend to the protection of execution transcripts. Consequently, data-dependent processing patterns such as memory traces and read/write volumes can still reveal critical information, risking privacy breaches (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016) even when the core data remains encrypted. To ensure strong privacy, modern SCA systems utilize data-oblivious primitives that exhaustively pad query processing complexities to a worst-case and data-independent upper bound (Bater et al., 2017; Liagouris et al., 2023; Poddar et al., 2021). However, such stringent protections can largely reduce system efficiency and hinder the generalization of conventional optimization techniques to SCA, which are typically data-dependent (Liagouris et al., 2023). To address this, recent efforts (Qin et al., 2022; Wang et al., 2021, 2022; Bater et al., 2018) have introduced Differentially Private SCA (DPSCA). This approach allows controlled information leakage under DP (Dwork et al., 2014) to mitigate constant worst-case overhead. For instance, systems under this model can dynamically compact an intermediate query size to a noisy estimate close to the actual size, avoiding exhaustive padding. As such, queries under DPSCA experience largely boosted efficiencies (e.g., up to $10^{5}\times$ faster (Wang et al., 2022)) compared to their “no leakage” counterparts. Despite these substantial performance gains, existing DPSCAs still face critical limitations that impede their practical uses, as elaborated below:
@@
$\bullet$ L-1. Unbounded privacy loss. Most DPSCA systems utilize a per-operator privacy expenditure model (Bater et al., 2018; Wang et al., 2021; Qin et al., 2022; Chu et al., 2021; Chan et al., 2022; Wang et al., 2022), meaning each query operator (e.g., join, filter) independently consumes a portion of the privacy budget. This approach can lead to either unbounded privacy loss or the forced cessation of query responses upon budget exhaustion. To mitigate this, some studies (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., 2023) propose private, locality-sensitive grouping, incurring a one-time privacy cost to pre-group data based on specific attributes. Subsequent queries on those attributes can be directly applied to a smaller subset and need no additional privacy budget. However, this method only supports simple queries (e.g., point and range); complex queries like joins still suffer from unbounded privacy loss.
@@
$\bullet$ L-2. Unoptimized execution plan. Conventional query planners can pre-estimate sizes for equivalent plans of a given query and select the most efficient plan with minimized intermediate sizes before execution (Silberschatz et al., 2011; Blasgen et al., 1981). In contrast, SCA systems lack this capability, and even DPSCA designs (Qin et al., 2022; Wang et al., 2022; Zhang et al., 2023; Wagh et al., 2021; Bater et al., 2018, 2020) can only reactively determine plan sizes during runtime. This inherent limitation often forces existing systems to settle for less efficient query plans, such as costly join orders, which lead to significantly inflated intermediate sizes (§ 7.2) and substantially hinder performance.
@@
$\bullet$ L-3. Lossy processing. Noise from randomized mechanisms in DPSCA also introduces a unique accuracy issue (e.g., conventional DP mechanisms may generate negative noise, applying which to obfuscate the sizes of intermediate query results can cause losing qualified real tuples), and unfortunately, no existing DPSCA can mitigate such loss for complex queries (Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Zhang et al., 2023; Groce et al., 2019). Furthermore, stronger DP settings can further increase noise variance, which amplifies errors, significantly impacting the utility of SCA systems.
1.1. Overview of SPECIAL
@@
In this work, we introduce SPECIAL, an innovative SCA system that resolves the aforementioned limitations all at once through a new paradigm called synopsis-assisted secure processing. At its core, SPECIAL incurs a one-time privacy cost to gather DP synopses (statistics of base tables) from owners’ data. These synopses are then used to accelerate complex query processing, and enhance SCA query planning. Notably, SPECIAL is the first system to provide all of the following benefits: (1) Bounded privacy—the privacy loss in SPECIAL is strictly limited to the one-time synopses release stage, with absotely no additional privacy cost during complex query processing and planning; (2) Advanced query planning—it builds an advanced SCA planner that can exploit plan sizes before runtime; and (3) Lossless processing—it ensures exact results with no data omissions. An overview of SPECIAL is shown in Figure 1.
@@
SPECIAL operates under a standard server-aided MPC model (Kamara et al., 2011) with three key participants: data owners, at least two SPECIAL servers, and a vetted analyst. The process begins with data owners securely outsourcing their data, typically through secret sharing (Kamara et al., 2011; Wang et al., 2021; Zhang et al., 2023; Wang et al., 2022), and privately releasing corresponding DP synopses (§ 4) to the servers. SPECIAL introduces a set of novel primitives (§ 5) that can leverage these synopses to accelerate secure query operations. Once the data and synopses are in place, analysts can submit Select-Project-Join-Aggregation (SPJA) queries (Silberschatz et al., 2011) for analytics. To process queries, a private planner (§ 6), running on SPECIAL servers, strategically orchestrates SPECIAL primitives to process the query and optimize performance. Finally, the results are securely returned to the analyst.
@@ 1.2. Unique challenges and key contributions
$\bullet$ C-1. How to select proper synopses? Even for a single relation, one can find numerous attribute combinations for generating synopses. Improper selection can lead to large errors (e.g. using too many synopses or high-dimensional attributes (Zhang et al., 2021)), or reduced functionalities (e.g., using only simple attributes (Zhang et al., 2023)). Hence, a key challenge is selecting a limited set of DP synopses to optimize the privacy budget for complex query processing. Our approach is informed by two observations: (i) secure joins are resource-intensive and need prioritized acceleration, and (ii) synopses for common filtering predicates are vital as they allow pre-built indexes on base relations for fast access. Consequently, we propose a focused strategy (§ 4) that targets low-dimensional (1D and 2D) attributes frequently involved in joins and filters within a representative workload.
@@
$\bullet$ C-2. How to enforce lossless processing? Private synopses do not immediately implies lossless guarantees. Thus, a second challenge is designing practical approaches to achieve lossless results without violating privacy goals. To address this, we employ one-sided DP noise (either strictly positive or negative, § 4) in generating synopses, and design novel primitives (§ 5) based on them to pessimistically estimate filter cardinalities and intervals of index structures. To ensure lossless processing of complex joins, we extend upon cutting-edge join upper bound techniques (Hertzschuch et al., 2021) to privately estimate lossless join sizes using DP synopses. To our knowledge, this is the first study to support private join upper bound estimation.
@@
$\bullet$ C-3. How DP synopses can empower efficient query processing? The use of DP synopses in SCA is largely underexplored, leaving a knowledge gap regarding their potential to enhance query efficiency. To navigate this potential, we explore various use of synopses in accelerating secure processing including private indexes SPEidx (§ 5.2), and compacted oblivious operations SPEop (§ 5.3). We also design a novel private query planner (§ 6) that efficiently orchestrate the execution of SPECIAL primitives (e.g., SPEidx, SPEop) to process SPJA queries. The planner uses available synopses to privately estimate intermediate result sizes and operation costs for a set of equivalent execution plans of a given query. It then executes the one with the lowest estimated cost.
@@
$\bullet$ C-4. How to systematically evaluate SPECIAL? A major is the absence of open benchmarks. We address this by initiating an open-source evaluation set, accessible to the public. Specifically, we use public financial data (Rel, [n.d.]) and design eight test queries, ranging from simple linear queries to complex 5-way joins. We also re-produce an open version of the HealthLNK benchmark. We evaluate our prototype, SPECIAL, against the state-of-the-art (SOTA) DPSCA system, Shrinkwrap (Bater et al., 2018), and the conventional SCA system, SMCQL (Bater et al., 2017). Results indicate that SPECIAL outperforms Shrinkwrap, reducing query latency by up to $80.3\times$, and SMCQL, with at least a $114\times$ reduction in query latency. Additionally, SPECIAL improves memory efficiency in complex join processing by more than $900\times$ compared to both systems. Moreover, scaling experiments show that SPECIAL can effectively scale up to 8.8M rows dataset and up to 9-way complex joins. All benchmarks, including our prototype implementation, are open-sourced and available at (Lovingmage, 2024).
@@ 2. Background
General notations. We consider the logical database $\mathcal{D}$ to contain multiple private (base) relations $\{D_{1},D_{2},$...\}, where each relation $D_{i}$ is owned by a specific party $P_{i}$. A base relation $D$ (we omit subscript for simplicity) has a set of attributes $attr(D)$. The domain of an attribute $A\in attr(D)$ is denoted by $dom(A)$, and the combined domain of a collection of attributes $\mathbf{A}=\{A_{1},A_{2},...\}\subseteq attr(D)$ is denoted by $dom(\mathbf{A})=\prod_{A\in\mathbf{A}}dom(A)$. For a tuple $t\in D$, and $\mathbf{A}\subseteq attr(D)$, we use $t.\mathbf{A}$ to denote the attribute value of $\mathbf{A}$ in $t$. A logical query, represented by $q(\mathcal{D})$, applies transformations and computations on $\mathcal{D}$. In this work, we focus on SPJA (Silberschatz et al., 2011) queries.
@@
Frequency (count). Given $D$, $\mathbf{A}\subseteq attr(D)$, and a set of values $\mathbf{v}\in dom(\mathbf{A})$, the frequency (count) of $\mathbf{v}$ in $D$ is the total number of tuples $t\in D$ with $t.\mathbf{A}=\mathbf{v}$. In addition, the max frequency moments (MF) of $\mathbf{A}$ is defined as ${\mathtt{mf}}(\mathbf{A},D)=\max_{\mathbf{v}\in dom(\mathbf{A})}|\{t\in D\mid t.\mathbf{A}=\mathbf{v}\}|$.
@@
Histograms. Given $D$, and $\mathbf{A}\subseteq attr(D)$, the (equal-width) histogram $\mathbf{h}(\mathbf{A},D)=(c_{1},c_{2},...,c_{m})$ is a list of counts for the attribute values in $\mathbf{A}$. Specifically, $\mathbf{h}$ partitions $dom(\mathbf{A})$ into $m$ “equal-sized” domain intervals $(B_{1},...,B_{m})$, and a count $c_{i}\in\mathbf{h}$ is the number of tuples $t\in D$ with $t.\mathbf{A}$ in the interval of $B_{i}$.
Query planning. Modern databases parse queries into physical plans (Silberschatz et al., 2011) that can be executed by the underlying query engine. These plans specify the operations like scans, joins, and sorts, and the order in which they’re performed. The same query can have multiple equivalent execution plans, but their performance can vary greatly depending on resource usage and data access patterns. Query planning (Blasgen et al., 1981), done before runtime, involves selecting cost-efficient plans from these options. A key part of this process is accurately estimating intermediate result sizes, known as cardinality estimation (CE) (Harmouch and Naumann, 2017), which relies heavily on table statistics. Two crucial statistics in modern cardinality estimation methods are: (i) histograms, which are vital for estimating selectivities in filters, and (ii) max frequency, which is crucial for estimating join sizes.
[…]
@@
Differential privacy (Dwork et al., 2014). DP ensures that modifying a single input tuple to a mechanism produces only a negligible change in its output. To elaborate, consider $D$ and $D^{\prime}$ as two relations differing by just one tuple, then DP defines the following.
Definition 2.1 ($(\epsilon,\delta)$-DP).